### v3

# Data Distributional Properties As Inductive Bias for Systematic Generalization

Abstract
Deep neural networks (DNNs) struggle at systematic generalization (SG). Several studies have evaluated the possibility of promoting SG through the proposal of novel architectures, loss functions, or training methodologies. Few studies, however, have focused on the role of training data properties in promoting SG. In this work, we investigate the impact of certain data distributional properties, as inductive biases for the SG ability of a multi-modal language model. To this end, we study three different properties. First, data diversity, instantiated as an increase in the possible values a latent property in the training distribution may take. Second, burstiness, where we probabilistically restrict the number of possible values of latent factors on particular inputs during training. Third, latent intervention, where a particular latent factor is altered randomly during training. We find that all three factors significantly enhance SG, with diversity contributing an 89% absolute increase in accuracy in the most affected property. Through a series of experiments, we test various hypotheses to understand why these properties promote SG. Finally, we find that Normalized Mutual Information (NMI) between latent attributes in the training distribution is strongly predictive of out-of-distribution generalization. We find that a mechanism by which lower NMI induces SG is in the geometry of representations. In particular, we find that NMI induces more parallelism in neural representations (i.e., input features coded in parallel neural vectors) of the model, a property related to the capacity of reasoning by analogy. Our code is available at: https://github.com/fdelrio89/data-systematic
1 Introduction
Humans excel at adapting quickly to novel tasks. This feature emerges from our systematic generalization (SG) ability, that is, the flexible recombination of previous (blocks of) knowledge in ways that allow us to behave effectively under novel task demands and, more generally, help us make sense of the world [23, 52].
Importantly, SG depends on prior knowledge or inductive biases upon which agents can rely on [46]. Indeed, inductive biases enable the representing of knowledge in the form of discrete independent concepts, which can be leveraged for flexible and rapid adaptation to new situations. In humans, for example, visual perception is strongly biased towards shapes which aids in recognizing previously unknown objects [28, 25]. This bias helps us reason independently about color, texture, or form, aiding in recognizing unknown objects regardless of their specific combination of properties. A child who has seen many horses, for instance, is likely to infer that a zebra moves similarly to a horse despite their differing colors.
In contrast to empirical observations in humans [13], SG does not emerge naturally in an i.i.d. setting in deep neural networks (DNNs) [43, 71, 39, 38]. Since training data is always finite in machine learning, models need built-in assumptions to generalize to novel inputs. These assumptions, often called inductive biases [27], help guide models toward solutions in situations not seen during training, and promote model weight configurations (i.e., solutions) that encourage generalization aligned with specific objectives [76, 8]. Inductive biases can take different forms, such as constraints over loss functions [41], architectures [51], pre-training [15], learning [45], or the geometry of representations [70, 72, 34]. However, the role of specific training data distributional properties as inductive biases remains surprisingly understudied.
A notable exception to this trend is the study of data complexity. Indeed, both in terms of patterns and scale, increasing training data complexity can improve SG in language models [44, 36, 62, 80, 1, 22]. Similarly, within the realm of vision, augmenting training data with translations has been shown to improve learned invariances [10]. Nevertheless, several questions remain open. First, although incrementing training data complexity does promote SG, the property (or properties) of data distribution that subtends this promotion remains obscure. Second, despite the tremendous success and broad applicability of vision-language models [49, 47, 7, 48], whether data training properties can boost SG in these models has not been explored. Third, a systematic analysis of the effect of training data properties on internal representations and how these can promote SG is also missing. In other words, we lack a functional explanation of how data properties may promote SG.
In this work, we focus on the following research questions: "Can SG be induced solely through changing properties of the training data? If so, how and why does this work?". To answer these questions, we focus on a Multi-modal Masked Language Modeling (MMLM) setting and test for three different properties of the data: diversity, an increase in the cardinality of latent factors of the training distribution; burstiness, a limitation on the cardinality of latent factors in particular instances of the training data; and latent intervention, a targeted intervention on the value of a latent factor. We construct datasets that alter these properties by manipulating their generative latent factors and test how a model trained on these datasets generalizes to instances of unseen combinations of latent factors. Our experiments demonstrate that manipulating these properties increases the SG capability of a model.
Finally, to understand why these properties enhance SG, we conduct thorough experiments. We study the role of the Normalized Mutual Information (NMI) between latent attributes of a dataset and find that it correlates strongly with out-of-distribution (OOD) generalization. We then turn to the effect of data distributional properties on learned representations. Moreover, we examine the representational geometry relative to baseline models. A prevailing conjecture, inspired by biological neural representations, suggests that the more parallel representations remain under a particular change in a latent attribute, the better they are at SG. Our experiments reveal that NMI influences this geometry. Thus, our contributions are summarized as follows:
  • We are the first to study the impact of training data properties and its effect on SG in the multimodal setting.
  • We show that increasing the number of values of a given data factor, a property of data we call diversity, promotes SG in an MMLM task. We achieve an absolute increase in accuracy of up to 89% in a systematic test set, solely through manipulation of the training data.
  • Analogously, we show that probabilistically restricting the number of values of latent factors in a given input, a term we call burstiness, also increases SG, with an absolute 15% increase over the baseline. This increase, however, comes at the cost of decreased performance related to the restricted factor.
  • We also show that intervening some of the latent factors in the training data, without violating systematicity, significantly increases SG, up to an absolute increase of 15%.
  • While it is known that lower NMI between latent factors tend to produce models with better SG performance, we find that for datasets with the same level of NMI, performance differs greatly when altering dataset diversity. This suggests a different mechanism by which diversity increases OOD generalization.
  • We find that one mechanism by which lower NMI induces better SG is by promoting the emergence of representations with more parallellism under changes in latent attributes. To the best of our knowledge, this mechanism for SG has not been previously reported.
2 Problem Formulation
We train a model on a similar task to CLEVR [37] to correctly respond to a textural query given contextual information (in our case, images) about a group of entities contained in it. In particular, queries that will give a partial description of the image (e.g., "small red rubber sphere, large [?] metallic cylinder, etc.") will always resolve to predict the value of one or more particular attributes (e.g., color in this case) of a particular entity (e.g., cube). See Figure 1 for an example input image, query, and expected answer. A model able to perform SG will achieve this over a test set with unseen combinations of these attributes during training. This is relevant because models tend to memorize combinations of attributes (which may prove spurious) during training rather than learn each one separately.
2.1 Task formalization
Let $A$ be a set of latent attributes $\{a_{1},a_{2},...,a_{|A|}\}$. Each attribute $a_{i}$ may take $|a_{i}|$ possible values. Let an entity $e$ be described by particular realizations of $A$. Thus, an entity $e_{i}$ can be written as an $|A|$-dimensional vector.
We define a context $c$ as a lossless representation of $N_{c}$ entities. Let $q$ be one or more queries about the value of a particular attribute of an entity in $c$, Let $y$ be the correct answer to queries in $q$.
Altogether, we define $\mathcal{D}=\{(c_{i},q_{i},y_{i}),i=1..N\}$ to be a dataset of $N$ samples, where $c_{i},q_{i},y_{i}$ refer to the context, queries and answers for the $i$-th element in the dataset. The task consists in predicting the correct $a_{i}$, given $c_{i}$ and $q_{i}$.
2.2 Evaluation
To perform our evaluation, as is standard in the systematic generalization literature, we will assume two latent attributes $\{a,b\in A\}$, where $b$ is the attribute we wish to predict systematically. We will work with two data distributions; Split A: where properties $a$ and $b$ are related; and Split B, where this relation does not hold. Specifically, Split A will contain a subset of all combinations of $a$ and $b$, while Split B will contain all remaining combinations. During our experiments, we use three different sets for training and evaluation: training ($\mathcal{D}_{train}$), test in-distribution ($\mathcal{D}_{test-ID}$), and test out-of-distribution ($\mathcal{D}_{test-OOD}$). $\mathcal{D}_{train}$ and $\mathcal{D}_{test-ID}$ come from Split A, while $\mathcal{D}_{test-OOD}$ comes from Split B. During evaluation we will focus on performance on property $b$.
2.3 Distributional Properties of Training Data
We wish to study how three properties of $\mathcal{D}_{train}$ affect SG. Therefore, we look for distributional properties of $\mathcal{D}_{train}$ that might be able to break the link between $a$ and $b$. To achieve this, we propose three properties, define them, and give intuition on why they may work to achieve this goal.
2.3.1 Diversity
To disrupt the link between properties $a$ and $b$, we propose to increase the total number of values $a$ may take. We hypothesize that increasing the diversity of values of $a$ on the training set might make it harder for the model to memorize all combinations of $a$ and $b$. Thus, the model may prefer solutions that learn to detect $a$ and $b$ separately. Therefore, we propose diversity over attribute $a$ as the total cardinality of $a\in A$, that is $div(a)=|a_{i}|$. Over our experiments, we modulate diversity over $a$ in a dataset by altering the number of possible values $a$ may take in the training set.
2.3.2 Burstiness
Another way to influence the link between $a$ and $b$ is to disrupt the learning of $a$, such that the model may not rely on $a$ to predict $b$. To instantiate this idea as a data property, we propose the use of burstiness. Burstiness over $a$ is defined as limiting the number of possible values $a$ may take within an individual context $c$. Intuitively, this makes batches of data where only some value of $a$ will be heavily featured, while on other batches a different set of values for $a$ will dominate. That is our intuition of burstiness. To modulate burstiness, we define a probability $p_{burst}$ that a sample is bursty or not. This property is adapted for use in the MMLM setting from previous work [11], where it was applied by limiting the number of different class labels and eliciting a meta-learning algorithm.
2.3.3 Latent Intervention
Finally, another way to break this link is to directly alter the value of $a$ while keeping the value of $b$ constant. This approach is reminiscent of the causal inference literature [64, 65], which uses interventions over causal variables to learn interventional probability distributions, that is, the probability distribution of outcomes should a certain intervention be affected. Thus, latent intervention of $a$ is related to altering the values of a latent attribute consistently across all entities in a context $c$. We modulate the degree of latent intervention by the amount the attribute is altered.
3 Experimental Setup
We proceed to detail how we operationalize the ideas from the previous section.
3.1 Dataset
We create a dataset resembling the structure of the CLEVR dataset [37] by leveraging their generative code. Attributes for this dataset are $A=\{shape,color,material,size\}$. Our main goal is to predict the shape attribute systematically, while excluding certain combinations of shape-color from $\mathcal{D}_{train}$. There are three possible shapes: {sphere, cube, cylinder}. The context for this dataset is a 224 x 224 color image of 3 to 10 objects. Queries are defined as a textual description of the scene where the property of the object that is to be predicted is masked. Contextual information is also included in the query to make queries unambiguous. Our training set consists of 75,000 samples and the test set of 15,000.
Properties $a$ and $b$ will be color and shape, respectively. We designate a subset of shapes (cubes and cylinders) to introduce systematic differences. That is, some combinations of these shapes and colors will be in Split A, while the rest will be in Split B. Spheres, on the other hand, will combine to all colors in every split, so they may act as a control.
3.2 Distributional Properties of the Training Data
We modulate diversity by altering the number of colors available in the $\mathcal{D}_{train}$. We test the following total number of colors: $\{8,27,64,125,216\}$. These values come from dividing the color span in each channel in $n\in\{2..6\}$ equal parts. Burstiness is implemented by limiting the number of possible colors in a given image to 3. This property is modulated by altering $p_{burst}$ with the following values: $\{0.0,0.5,1.0\}$. Finally, we implement latent intervention by randomly altering the hue of all colors in the image via the ColorJitter transform in PyTorch [61]. Our experiments use the following values: $\{0.0,0.05,0.1,0.5\}$.
3.3 Model
We use a Transformer encoder [75] that receives both the context and query as inputs. Following standard procedure [16], images are divided into patches and arranged into a sequence of embeddings via a linear layer. The text is passed through an embedding layer. Learned positional and modality (image or text) embeddings are added to the input before the encoder. Both sequence embeddings are concatenated and fed into the transformer encoder (see Fig. 1).
3.4 Evaluation
We run each experiment for 3 seeds and report their mean and standard error. All metrics are related to performance on cubes and cylinders to correctly assess generalization in objects with systematic differences to the training set.
Figure 1: (A) Example input for our task. Image shows objects with different properties, while query gives a textual description of the objects while masking properties to be predicted. Sample query lists numbered objects from 1-3. (B) Diagram of the model during training. Context is an image, while the query is text describing the scene, with some text related to properties of objects being masked. Context and query enter into a Transformer that outputs its predictions for the values of the masked properties.
4 Baseline Results on SG
We first analyze the behavior of a baseline model trained on 8-colors in $\mathcal{D}_{train}$, probing its ability to predict specific attributes. The results are shown in Fig. 2.
Figure 2: Accuracy of the baseline model (8 colors) for predicting different properties across data splits. While ID performance remains high, OOD performance drops sharply for shape and color, indicating the model learns shape-color combinations as features, rather than achieving SG. Surprisingly, material and size also show a drop in OOD performance, even though the model has been exposed to all combinations of these attributes in training.
ID performance remains similar to that of $\mathcal{D}_{train}$ for all tasks. However, delving into performance for $\mathcal{D}_{test-OOD}$ reveals three distinct generalization patterns. First, unexpectedly, the material and size show a decrease in $\mathcal{D}_{test-OOD}$ performance, even though the model has been exposed to all combinations of these attributes during training. Second, we observe a significant decrease in performance for the color task (44% drop) and shape (99% drop). This is especially bad for the shape task which achieves much worse than random performance (33% to predict between three shapes in $\mathcal{D}_{test-OOD}$). These results strongly suggest that the model faces challenges when performing SG. These results confirm previous findings that indicate DNNs encounter significant performance challenges when required to generalize systematically [43, 71, 39, 38]. Third, given that both shape and color are related systematically, a natural question arises: why is color less affected than shape? An explanation for this difference in performance between color and shape on $\mathcal{D}_{test-OOD}$ emerges from how image information is fed to the model. Due to the division of images into patches, the model can predict the color of a particular object solely based on the information of one patch. However, to predict the shape of an object, the model needs information from additional patches, making this ability harder to train and thus successfully achieve SG.
We turn to analyzing how modulating the amount of training data affects SG. A popular way of promoting generalization is to increase the amount of training data. Therefore, we train models on different percentages of $\mathcal{D}_{train}$. Results are shown in Figure 3(a). While ID roughly stays the same, increasing the amount of training data does not result in better OOD generalization. Performance degrades as dataset size increases. This suggests that merely adding more ID training data may not be a robust way of inducing a bias toward SG. Even worse, it seems to double down on whatever non-systematic bias the model is learning. We now alter the distribution of $\mathcal{D}_{train}$.
(a) Shape task (8 Colors)
(b) Shape task (216 Colors)
Figure 3: Accuracy on the shape task with varying training data sizes for models trained on 8 colors (baseline) vs. 216 colors. (a) Increasing dataset size does not improve OOD performance and slightly degrades it. (b) Greater diversity significantly enhances OOD generalization, with models trained on just a quarter of the data severely outperforming the 8-color baseline.
5 Altering the Training Distribution for SG
We investigate the impact on SG of altering $\mathcal{D}_{train}$ by modifying properties that disrupt the link between $a$ and $b$.
5.1 Diversity
(a) Color Task
(b) Shape Task
(c) Material Task
(d) Size Task
Figure 4: Accuracy versus the number of colors in $\mathcal{D}_{train}$ for $\mathcal{D}_{test-ID}$ and $\mathcal{D}_{test-OOD}$ for the color and shape tasks. Performance for the shape task increases drastically for the OOD split as we increase colors, increasing 86% in absolute terms over the 8-color baseline. Moreover, performance in the color task also tends to increase in the OOD split, while ID only suffers slightly, even though the task becomes significantly harder. Remarkably, the material and size task rapidly increase their $\mathcal{D}_{test-OOD}$ performance as color increases.
To increase diversity, we run experiments in which we increase the colors in $\mathcal{D}_{train}$ from 8 to 216. As shown in Figure 4, greater color diversity improves SG on the shape task (see the monotonic increase of green bars). An absolute improvement of $89\%$, a staggering amount that reflects the difference between a model that does not generalize systematically and one that does. Surprisingly, the test-ID color prediction drops only slightly ($\sim 6\%$) (Fig. 4(a)), despite the increased difficulty of an increased class count. Moreover, higher color diversity reduces the performance gap between $\mathcal{D}_{test-ID}$ and $\mathcal{D}_{test-OOD}$ in the color task (Fig. 4(a)). Remarkably, OOD performance for the $material$ and $size$ tasks also increases significantly when adding colors. This suggests that the increase in diversity appears to be creating an inductive bias where at least color is being disassociated from shape, material, and size. Moreover, this increase for material and size occurs rapidly with just 27 colors, while for shape, this keeps increasing until reaching 216 colors.
Finally, we assess whether increased diversity affects sample efficiency. We train models on datasets with 216 colors of $\{0.25,0.5,0.75\}$ the original training size and compare them to the 8-color baseline at full size. The results are shown in Figure 3(b). Remarkably, even with just $25\%$ of the data with increased diversity, produces much stronger OOD generalization than the baseline, while the gap between ID and OOD remains relatively low.
5.2 Burstiness
(a) Color Task
(b) Shape Task
Figure 5: Change of accuracy with respect to no burstiness for the color and shape tasks in test-ID and test-OOD for different levels of burstiness over color for various numbers of colors. Limiting the number of colors available for each image during training allows the model to gain up to 14.8% more accuracy over the baseline. The color task, however, suffers up to 14.3% decline as the color task becomes easier to memorize.
As was mentioned in Section 2.3, another way of disrupting the link between color and shape, is to hinder the learning of color by limiting the number of colors contained in a given image, something we call burstiness. Figure 5 shows the results. As can be seen, increasing the burstiness of the training data increases the SG capability of the model in the shape task by up to 14.8%. As expected, however, the color task is significantly affected, both in $\mathcal{D}_{test-ID}$ and $\mathcal{D}_{test-OOD}$, losing up to 14.3% accuracy in $\mathcal{D}_{test-OOD}$ for 64 colors. This last effect is especially pronounced as we increase the number of colors contained in the dataset, which is to be expected as the color task becomes significantly harder as the number of colors increases dramatically, while the limit of only 3 colors per image remains fixed. Finally, the effect of burstiness on material and size is very small, with at most 1% gain in $\mathcal{D}_{test-OOD}$.
5.3 Latent Intervention
Figure 6 shows results for modulating latent intervention over color. In our experiments, we can observe that increasing the degree of the perturbation we apply to the color of the images during training increases the model performance on the shape task by up to 15%, a non-trivial amount while retaining OOD performance on the color task. For the material and size tasks, latent intervention adds up to 2% in $\mathcal{D}_{test-OOD}$ performance, and the effect is reduced as the number of colors increases.
(a) Color Task
(b) Shape Task
Figure 6: Change in accuracy after applying latent intervention for the color and shape tasks in test-ID and test-OOD for different levels of latent intervention of the color latent attribute for various numbers of colors. X-Axis not at scale. Altering the color hue randomly during training allows the model to gain up to 15% more OOD accuracy over baseline.
5.4 Discussion
While all properties improve SG in the shape task, diversity clearly seems to have the greatest impact overall. However, we see that the other properties seem to be complementary to diversity. This suggests a positive answer to our research question, “Can SG be induced solely through changing properties of the training data?".
Surprisingly, altering these properties affects not only performance on $\mathcal{D}_{test-OOD}$ for shape and color, but also material and size. This suggests that, by default, a model tends to entangle all latent features in its representation. We demonstrate that modifying the training distribution to specifically target spurious relations between latent factors can alter this behavior. We explored three approaches with varying success, addressing our second question: “How can we modify the training distribution to induce SG?". Finally, we turn to understanding why these properties influence SG, beginning with an analysis of model capacity.
6 Lack of Capacity does not explain SG
Given that diversity has shown to improve SG performance, we would like to gain a deeper understanding of why this is the case. Our first hypothesis is the following: adding more colors to the training distribution pressures the capacity of the model, forcing it to learn simpler solutions where shape and color are represented independently. In other words, lack of capacity may force the model to find a more systematic solution. To test this, we test two cases: (1) We artificially restrict the capacity of a baseline Transformer model trained on 8 colors by limiting the size of the residual stream dimension ($d\in\{32,64,128\}$). (2) We augment the capacity of the model trained on 216 colors by increasing this dimension by $\{512,1024\}$.
Figure 7 shows results for these experiments. As can be seen, performance in both $\mathcal{D}_{test-ID}$ and $\mathcal{D}_{test-OOD}$ decreases as model capacity is reduced in the first case, and the converse happens when we increase capacity. So, lack of capacity does not seem to be the mechanism by which diversity achieves SG. We now turn to mutual information between latent attributes as a possible explanation.
(a) Color Task
(b) Shape Task
(c) Color Task
(d) Shape Task
Figure 7: ID and OOD accuracy for models trained with varying hidden dimensions on a 8 (a-b) and 216 (c-d) color dataset for the color and shape task. (a-b) With low diversity (8 colors) both metrics drop as capacity decreases. (c-d) With higher diversity, both slightly improve as capacity increases. These suggest that a capacity-bottleneck is not the cause for achieving greater SG when augmenting the number of colors in the training set.
7 Mutual Information versus Diversity
Figure 8: NMI in the training set vs OOD accuracy for the shape task for datasets with 8, 64 and 216 colors. For lower NMI, there is greater OOD performance. However, for similar levels of NMI but differing levels of diversity, we observe that greater diversity increases performance. This suggests that the effect of diversity is not only related to changing the NMI in the dataset.
Previous work [1] links Normalized Mutual Information (NMI) of latent attributes and SG. Mutual information quantifies the amount of information one variable (e.g., color) provides about another (e.g., shape). Then, this value is normalized with the entropy of the distributions to produce the NMI. The closer NMI is to zero, the more independent these factors are. In other words, a near-zero mutual information score indicates that the dataset offers strong evidence to the model that shape and color should be treated as independent attributes.
Changing diversity in a dataset impacts its NMI, so we proceed to study this effect. We seek to decouple the effect of increased diversity from a dataset’s NMI. To modulate NMI for a given level of diversity, we produce datasets that assign colors that both cubes and cylinders see during training (common colors) and a set of colors that are only seen either by cubes or by cylinders (exclusive colors). Modulating the ratio of common colors within the dataset directly alters its NMI, while keeping the number of colors constant.
To measure the NMI, we follow previous work [1] and focus on the NMI between the color and shape of all objects of the training dataset. It shows a Pearson correlation of -0.79. Figure 8 plots results for datasets with differing levels of diversity. For lower NMI, there is greater OOD performance. However, for similar levels of NMI but differing levels of diversity, we observe that greater diversity increases performance. This suggests that diversity influences more than just NMI, pointing to an additional mechanism. We hypothesize that this mechanism affects the geometry of representations, which we explore next.
8 Geometry of Representations
8.1 Disentanglement
The role of disentaglement in promoting SG is unclear. Some studies argue that disentanglement promotes SG by enforcing factorized representations that align with the underlying structure of the data [32, 33, 18]. While other research [56, 77, 73] finds no relation between disentanglement and OOD generalization. Using DCI metrics [20], we assess measures of Disentanglement and Completeness. Disentanglement quantifies how many latent attributes each representation dimension predicts, ranging from 0 (equal predictive power across all attributes) to 1 (each dimension exclusively predicts a single attribute). Completeness measures how distributed the prediction of each attribute is across dimensions, with 0 indicating equal contribution from all dimensions and 1 meaning a single dimension predicts each attribute.
Figure 9 shows results for these metrics for representations for color and shape task, from models trained with varying color diversity. Higher diversity increases disentanglement and completeness for shape task representations, while the converse is true for color task representations.
(a) Shape Task Vectors
(b) Color Task Vectors
Figure 9: Disentanglement and Completeness for models trained on different levels of color diversity for vectors predicting the shape and color tasks. Both metrics increase for the shape task vectors, while the contrary happens for color task vectors.
We also complement our analysis by drawing from neuroscience-inspired studies highlighting how parallelism in representations enhances generalization [9, 35].
8.2 Parallelism in Representations
The intuition behind studying parallelism in representations is that two objects with a single different factor should have a constant direction offset in their representation, independent of other factors or context. Thus, that factor is encoded in representational space as the direction of the difference between both objects’ representations.
Consider the classic example for analogy in DNNs: $king-man+woman=queen$ which implies: $king-man=queen-woman=royalness$.
Here, subtracting gender produces vectors that move in the same direction (an abstract property one might consider royalness) for two otherwise different vectors. Encoding this property consistently in a single direction allows the model to generalize that concept through analogy—a powerful inductive bias for solving new tasks [55].
We proceed to apply this same principle to the latent factors contained in our dataset. From a starting set of 1,024 images, we sample 3,500 pairs of representations from objects that share a common latent attribute. To represent each object we use the representation vector associated to the attribute we wish to study, which in this case will be either shape or color. We proceed to find for each of these pairs another pair of representations which represents a change in this attribute (i.e. for shape, from a pair of red and green cubes, we look for a pair of red and green cylinders). We substract these pairs and from the resulting pair we obtain a single value called a p-score [35]. This score is calculated by measuring the cosine of the angle between vectors in the resulting pair. We report the mean of this value over the 3,500 pairs averaged over 5 runs.
Figure 10: Computation of the p-score metric for vectors $V_{A}$ and $V_{B}$, representing the effect of ablating features on shape embeddings. We first subtract the embeddings to obtain vectors $V_{A}$ and $V_{B}$ and then calculate the degree of parallelism between them using cosine similarity. A high p-score indicates strong parallelism, suggesting that the attribute (in this case, cubeness) is encoded consistently in representational space.
We relate the p-score with performance on $\mathcal{D}_{test-OOD}$ for both the shape and color task for all models mentioned in Sections 4 and 5. Results for experiments with $8$, $64$, and $216$ colors are shown in Figure 11(a). Relating to the generalization capabilities in the shape task, we see a positive trend in OOD generalization when the p-score increases.
To assess how significant this relationship is, we compute the Pearson correlation between p-score and $\mathcal{D}_{test-OOD}$ performance on the shape task (details in the supplementary material). The result, $0.73$ with very low p-value, indicates a strong link between p-score and SG.
We also plot the relationship between NMI and p-score, as the positive relation both values have OOD generalization might be mediated by one or the other. Results are shown in Figure 11(b). A clear relationship between NMI and p-score appears: lower NMI datasets are related to higher p-scores in representations. Given that NMI is a stable property of the dataset, this suggests that lower NMI datasets tend to produce models with representations with higher p-scores. This suggests a novel mechanism by which lower NMI datasets produce inductive biases during learning that allow models to achieve greater OOD generalization. These results provide evidence that the dataset distribution influences the representational space geometry of the model biasing it towards OOD generalization.
(a) OOD accuracy vs. p-score
(b) p-score vs. NMI
Figure 11: (a) OOD accuracy on the shape task vs. p-score for models trained on 8, 64, and 216 colors. Showing a trend where higher parallelism scores in the representation correlate with better performance. (b) NMI vs p-score across models. Lower NMI datasets produce models with representations with higher p-score suggesting a mechanism by which lower NMI aids SG.
9 Related Work
Systematic Generalization.
Many works have addressed the problem of SG, proposing a variety of strategies. These include data augmentation methods [5, 3, 78, 2], training methods [44, 12, 45], adjustments to the loss function [30], architectural changes [14, 6, 19, 42, 24, 59, 67] and neuro-symbolic approaches [50, 57]. Currently, state-of-the-art methods are based on enhanced LLMs [17, 68].
Data-driven Inductive Biases.
Previous work has investigated the invariances [10], inductive biases such as shape bias [25], and robustness [22] acquired by deep learning models when trained on large-scale datasets. Prystawski et al. [66] shows the need for local structure in the training data for models to learn probabilistic reasoning abilities.
Other studies have examined the data properties that enable models to generalize OOD. Chan et al. [11] show that burstiness, long-tailed distributions and label mappings can foster in-context learning in Transformers. By contrast, our work focuses specifically on data properties that support SG. For SG, studies have shown that increasing the number of primitives or input length can enhance SG [63, 80, 69]. However, their focus is on generalization in the text domain, while we focus on images, examining how different latent factors within the dataset interact to drive SG.
Other work on generalization in the visual domain [1] has investigated SG in CLIP models, showing that statistical independence between object-attribute pairs and disentangled representations enhances SG. In contrast, we analyze the effects of explicit generative factors on SG and the mechanisms by which it is enforced.
Linear Representation Hypothesis
The parallelism of neural network representations is closely related to the linear representation hypothesis. This hypothesis suggests that neural network activations can be understood in terms of distinct features, which correspond to specific directions in activation space. Prior research has provided evidence for this hypothesis in both categorical [54, 58, 21, 53, 74, 4, 26, 60] and numerical features [29, 31]. Furthermore, it has been proposed that learning linear representations can enhance OOD generalization by reducing the nonlinear entanglement or contamination of different features [79].
10 Conclusion
We propose a different way of shaping inductive biases for SG, focusing on altering the properties of training data over other techniques. We show three properties of the training data that, when altered, significantly boost SG: diversity, burstiness, and latent intervention. Diversity has the strongest impact, improving SG by up to an absolute 89% over baseline, while the other factors complement it.
We also study why these are effective. We show that limiting model capacity is not a factor for SG, while increased diversity and lower NMI between latent attributes in the dataset seem to explain most of the gains in SG. Crucially, we test for a property of brain-like representations that measures parallelism in representations under changes in a latent attribute. We find that datasets with lower NMI tend to induce more of this parallelism, a property that allows models to extrapolate by analogy. This is, to the best of our knowledge, an unreported mechanism by which dataset distribution induces SG.
Acknowledgements
This work was supported by FONDECYT under Grants 1221425 and 11230762; and National Center for Artificial Intelligence (CENIA) under Grant FB210017 Basal ANID. We gratefully acknowledge the support of the AMD University Program (AUP), whose computing resources enabled the experiments conducted in this work.
Supplementary Material
11 Hyper Parameters
Images were resized to 224 × 224 pixels, and divided into patches of 16 × 16 pixels before being fed to the model. In the text, a masked language modeling (MLM) probability of 0.15 was used to randomly mask text tokens before feeding them to the model.
The model used during our experiments is a Transformer [75] with a hidden layer dimension of 256, with 4 transformer layers and 4 attention heads. Training was carried out using the Adam optimizer [40] with a learning rate of $1\times 10^{-4}$, a batch size of 256 and for 1000 epochs.
12 Dataset Samples
Below we display samples composed of images and their corresponding textual descriptions from the datasets described in Section 3.1:
Sample 1
Textual Description
small #00ff80 rubber cube [SEP] large #ffff80 rubber cylinder [SEP] large #80ffff rubber cylinder [SEP] small #ffff80 metal cube [SEP] large #000000 rubber sphere [SEP] small #80ffff rubber sphere [SEP] large #80ff00 rubber cylinder
Sample 2
Textual Description
large #005500 rubber cylinder [SEP] large #$5$aa metal cube [SEP] large #ff55aa rubber cube [SEP] small #ffff55 rubber sphere [SEP] large #aa00ff rubber sphere [SEP] large #00ff55 metal sphere
Sample 3
Textual Description
small #bf0040 rubber cylinder [SEP] large #40ff80 metal sphere [SEP] large #8040ff rubber cube [SEP] large #800000 rubber cube [SEP] small #bf4080 rubber cube [SEP] small #ffbfbf metal cylinder [SEP] small #ff4000 rubber cube [SEP] small #40bfff rubber cube [SEP] small #bfbf40 metal cube [SEP] large #ff8080 rubber sphere
13 Correlation between P-Score and OOD Performance
In Table 1 we show the Pearson correlation between different p-scores obtained for a model and their OOD performance on the shape task. Parallelism in the representations seems to strongly correlate with better OOD performance, suggesting that parallelism may be the property that is inducing SG in the models.
Metric Avg DIV BURST LI
Corr $0.73$ $0.76$ $0.69$ $0.73$
p-value $1.88e{-9}$ $7.17e{-4}$ $5.5e{-3}$ $3.45e{-5}$
Table 1: Pearson Correlation between P-Scores and OOD performance on the shape task broken down per training data property. DIV: Diversity, BURST: Burstiness, LI: Latent Intervention.
14 Detailed Experimental Results
In tables 2, 3, 4, 5, 6, and 7, we exposed the values obtained that were used to generate most of the figures in the work above. Table 2 shows the data used for Figure 2. Table 3 for Figure 3. Table 4 for Figure 4. Table 5 for Figure 6. Table 6 for Figure 5. And, Table 7 for Figure 7.
Train Test-ID Test-OOD
Color 94.3 ± 0.0% 94.1 ± 0.1% 50.5 ± 1.1%
Shapes 100.0 ± 0.0% 99.6 ± 0.0% 0.6 ± 0.1%
Materials 98.0 ± 0.0% 96.9 ± 0.0% 87.8 ± 0.9%
Size 98.1 ± 0.0% 98.0 ± 0.0% 91.2 ± 1.0%
Table 2: Table displaying values to form Figure 2. Accuracy for predicting different properties for different data splits for the baseline model (8 colors). ID performance for all tasks remains high, however OOD performance plummets for shape and color, suggesting the model is learning combinations of shape-color as features, instead of achieving SG. Unexpectedly, material and size also show a drop in OOD performance, even though the model has been exposed to all combinations of these attributes in training.
Color Shape
Train Fract. Test-ID Test-OOD Test-ID Test-OOD
8 Colors
1/8 91.5 ± 0.1% 47.4 ± 0.4% 97.1 ± 0.1% 2.1 ± 0.1%
1/4 93.1 ± 0.1% 47.4 ± 0.5% 99.0 ± 0.0% 1.2 ± 0.1%
1/2 93.9 ± 0.0% 47.3 ± 0.6% 99.2 ± 0.0% 0.6 ± 0.1%
1/1 94.1 ± 0.1% 50.5 ± 1.1% 99.6 ± 0.0% 0.6 ± 0.1%
216 Colors
1/8 22.8 ± 2.4% 5.0 ± 0.9% 45.1 ± 0.6% 43.3 ± 0.5%
1/4 79.6 ± 0.6% 60.4 ± 1.3% 89.1 ± 0.5% 81.0 ± 0.8%
1/2 86.1 ± 0.0% 74.1 ± 0.4% 94.4 ± 0.1% 88.4 ± 0.2%
1/1 88.1 ± 0.1% 77.9 ± 0.4% 96.3 ± 0.0% 90.0 ± 0.4%
Table 3: Table displaying values to form Figure 3. Accuracy for the shape task for different amounts of training data for the baseline model (8 colors) vs model trained on 216 colors. (a) Increasing dataset size does not increase OOD performance but rather degrades it slightly. (b) With increased diversity much stronger OOD generalization is achieved, with models trained on only a quarter of the data severely outperforming the 8-color baseline.
Color Shape Size Material
Num. Colors Test-ID Test-OOD Test-ID Test-OOD Test-ID Test-OOD Test-ID Test-OOD
8 94.1 ± 0.1% 50.5 ± 1.1% 99.6 ± 0.0% 0.6 ± 0.1% 98.0 ± 0.0% 91.2 ± 1.0% 96.9 ± 0.0% 87.8 ± 0.9%
27 91.7 ± 0.0% 67.5 ± 0.7% 96.9 ± 0.1% 1.5 ± 0.1% 99.5 ± 0.0% 97.6 ± 0.0% 98.6 ± 0.0% 96.7 ± 0.1%
64 90.8 ± 0.1% 73.0 ± 0.5% 96.9 ± 0.0% 48.5 ± 0.6% 99.7 ± 0.0% 97.6 ± 0.0% 99.2 ± 0.0% 96.8 ± 0.0%
125 89.0 ± 0.1% 67.0 ± 0.3% 96.1 ± 0.1% 81.8 ± 0.6% 99.8 ± 0.0% 97.8 ± 0.0% 99.2 ± 0.0% 96.9 ± 0.0%
216 88.1 ± 0.1% 77.9 ± 0.4% 96.3 ± 0.0% 90.0 ± 0.4% 99.8 ± 0.0% 97.7 ± 0.0% 99.3 ± 0.0% 97.2 ± 0.0%
Table 4: Table displaying values to form Figure 4. Accuracy versus the number of colors in $\mathcal{D}_{train}$ for $\mathcal{D}_{test-ID}$ and $\mathcal{D}_{test-OOD}$ for all tasks. Performance for the shape task increases drastically for the OOD split as we increase colors, increasing 86% in absolute terms over the 8-color baseline. Moreover, performance in the color task also tends to increase in the OOD split, while ID only suffers slightly, even though the task becomes significantly harder. Remarkably, the material and size task rapidly increase their $\mathcal{D}_{test-OOD}$ performance as color increases.
Color Shape Size Material
Num. Colors P Bursty Test-ID Test-OOD Test-ID Test-OOD Test-ID Test-OOD Test-ID Test-OOD
8 0.0 94.1 ± 0.1% 50.5 ± 1.1% 99.6 ± 0.0% 0.6 ± 0.1% 98.0 ± 0.0% 91.2 ± 1.0% 96.9 ± 0.0% 87.8 ± 0.9%
0.5 94.2 ± 0.0% 48.9 ± 0.9% 99.6 ± 0.0% 0.9 ± 0.2% 99.5 ± 0.0% 97.6 ± 0.0% 98.6 ± 0.0% 96.7 ± 0.1%
1.0 93.8 ± 0.0% 47.3 ± 0.1% 99.4 ± 0.0% 0.5 ± 0.0% 99.7 ± 0.0% 97.6 ± 0.0% 99.2 ± 0.0% 96.8 ± 0.0%
27 0.0 91.7 ± 0.0% 67.5 ± 0.7% 96.9 ± 0.1% 1.5 ± 0.1% 98.0 ± 0.0% 91.2 ± 1.0% 96.9 ± 0.0% 87.8 ± 0.9%
0.5 91.7 ± 0.0% 71.9 ± 0.2% 97.1 ± 0.0% 1.8 ± 0.3% 99.5 ± 0.0% 97.6 ± 0.0% 98.6 ± 0.0% 96.7 ± 0.1%
1.0 90.3 ± 0.0% 65.2 ± 0.8% 96.9 ± 0.0% 4.7 ± 0.4% 99.7 ± 0.0% 97.6 ± 0.0% 99.2 ± 0.0% 96.8 ± 0.0%
64 0.0 90.8 ± 0.1% 73.0 ± 0.5% 96.9 ± 0.0% 48.5 ± 0.6% 98.0 ± 0.0% 91.2 ± 1.0% 96.9 ± 0.0% 87.8 ± 0.9%
0.5 89.8 ± 0.5% 73.8 ± 1.0% 97.0 ± 0.1% 60.5 ± 2.0% 99.5 ± 0.0% 97.6 ± 0.0% 98.6 ± 0.0% 96.7 ± 0.1%
1.0 83.5 ± 0.2% 58.7 ± 1.0% 96.9 ± 0.1% 63.3 ± 0.9% 99.7 ± 0.0% 97.6 ± 0.0% 99.2 ± 0.0% 96.8 ± 0.0%
125 0.0 89.0 ± 0.1% 67.0 ± 0.3% 96.1 ± 0.1% 81.8 ± 0.6% 98.0 ± 0.0% 91.2 ± 1.0% 96.9 ± 0.0% 87.8 ± 0.9%
0.5 88.9 ± 0.1% 68.5 ± 0.3% 96.8 ± 0.0% 80.6 ± 1.2% 99.5 ± 0.0% 97.6 ± 0.0% 98.6 ± 0.0% 96.7 ± 0.1%
1.0 80.7 ± 0.2% 64.6 ± 0.3% 96.8 ± 0.0% 79.4 ± 0.5% 99.7 ± 0.0% 97.6 ± 0.0% 99.2 ± 0.0% 96.8 ± 0.0%
216 0.0 88.1 ± 0.1% 77.9 ± 0.4% 96.3 ± 0.0% 90.0 ± 0.4% 98.0 ± 0.0% 91.2 ± 1.0% 96.9 ± 0.0% 87.8 ± 0.9%
0.5 87.3 ± 0.0% 77.7 ± 0.2% 96.7 ± 0.0% 91.4 ± 0.2% 99.5 ± 0.0% 97.6 ± 0.0% 98.6 ± 0.0% 96.7 ± 0.1%
1.0 71.9 ± 0.2% 57.3 ± 0.4% 96.7 ± 0.0% 92.4 ± 0.3% 99.7 ± 0.0% 97.6 ± 0.0% 99.2 ± 0.0% 96.8 ± 0.0%
Table 5: Table displaying values to form Figure 5. Accuracy fo all tasks in test-ID and test-OOD for different levels of burstiness over color for various numbers of colors. Limiting the number of colors available for each image during training allows the model to gain up to 14.8% more accuracy over the baseline. The color task, however, suffers up to 14.3% decline as the color task becomes easier to memorize.
Color Shape Size Material
Num. Colors Jitter Test-ID Test-OOD Test-ID Test-OOD Test-ID Test-OOD Test-ID Test-OOD
8 0.00 94.1 ± 0.1% 50.5 ± 1.1% 99.6 ± 0.0% 0.6 ± 0.1% 98.0 ± 0.0% 91.2 ± 1.0% 96.9 ± 0.0% 87.8 ± 0.9%
0.05 94.1 ± 0.1% 46.6 ± 0.4% 99.6 ± 0.0% 0.4 ± 0.0% 99.5 ± 0.0% 97.6 ± 0.0% 98.6 ± 0.0% 96.7 ± 0.1%
0.10 94.1 ± 0.1% 48.0 ± 0.7% 99.6 ± 0.0% 0.9 ± 0.2% 99.7 ± 0.0% 97.6 ± 0.0% 99.2 ± 0.0% 96.8 ± 0.0%
0.50 93.9 ± 0.0% 47.1 ± 0.4% 99.7 ± 0.0% 0.3 ± 0.1% 99.8 ± 0.0% 97.8 ± 0.0% 99.2 ± 0.0% 96.9 ± 0.0%
27 0.00 91.7 ± 0.0% 67.5 ± 0.7% 96.9 ± 0.1% 1.5 ± 0.1% 98.0 ± 0.0% 91.2 ± 1.0% 96.9 ± 0.0% 87.8 ± 0.9%
0.05 91.9 ± 0.0% 69.4 ± 1.5% 97.4 ± 0.0% 1.4 ± 0.2% 99.5 ± 0.0% 97.6 ± 0.0% 98.6 ± 0.0% 96.7 ± 0.1%
0.10 91.8 ± 0.1% 68.4 ± 1.0% 97.3 ± 0.1% 1.7 ± 0.2% 99.7 ± 0.0% 97.6 ± 0.0% 99.2 ± 0.0% 96.8 ± 0.0%
0.50 91.9 ± 0.0% 66.1 ± 2.0% 97.2 ± 0.0% 1.1 ± 0.2% 99.8 ± 0.0% 97.8 ± 0.0% 99.2 ± 0.0% 96.9 ± 0.0%
64 0.00 90.8 ± 0.1% 73.0 ± 0.5% 96.9 ± 0.0% 48.5 ± 0.6% 98.0 ± 0.0% 91.2 ± 1.0% 96.9 ± 0.0% 87.8 ± 0.9%
0.05 90.9 ± 0.0% 68.5 ± 0.6% 97.1 ± 0.0% 57.1 ± 0.1% 99.5 ± 0.0% 97.6 ± 0.0% 98.6 ± 0.0% 96.7 ± 0.1%
0.10 91.0 ± 0.0% 67.1 ± 0.3% 97.3 ± 0.0% 62.3 ± 3.0% 99.7 ± 0.0% 97.6 ± 0.0% 99.2 ± 0.0% 96.8 ± 0.0%
0.50 91.1 ± 0.0% 68.6 ± 0.6% 97.8 ± 0.0% 63.8 ± 0.9% 99.8 ± 0.0% 97.8 ± 0.0% 99.2 ± 0.0% 96.9 ± 0.0%
125 0.00 89.0 ± 0.1% 67.0 ± 0.3% 96.1 ± 0.1% 81.8 ± 0.6% 98.0 ± 0.0% 91.2 ± 1.0% 96.9 ± 0.0% 87.8 ± 0.9%
0.05 89.8 ± 0.0% 71.1 ± 0.4% 96.8 ± 0.1% 85.0 ± 0.5% 99.5 ± 0.0% 97.6 ± 0.0% 98.6 ± 0.0% 96.7 ± 0.1%
0.10 89.7 ± 0.1% 68.4 ± 0.5% 96.7 ± 0.1% 79.2 ± 0.7% 99.7 ± 0.0% 97.6 ± 0.0% 99.2 ± 0.0% 96.8 ± 0.0%
0.50 89.8 ± 0.1% 70.6 ± 1.1% 96.9 ± 0.0% 85.8 ± 0.9% 99.8 ± 0.0% 97.8 ± 0.0% 99.2 ± 0.0% 96.9 ± 0.0%
216 0.00 88.1 ± 0.1% 77.9 ± 0.4% 96.3 ± 0.0% 90.0 ± 0.4% 98.0 ± 0.0% 91.2 ± 1.0% 96.9 ± 0.0% 87.8 ± 0.9%
0.05 88.4 ± 0.0% 79.8 ± 0.3% 96.8 ± 0.1% 92.2 ± 0.1% 99.5 ± 0.0% 97.6 ± 0.0% 98.6 ± 0.0% 96.7 ± 0.1%
0.10 88.0 ± 0.1% 75.6 ± 0.2% 96.4 ± 0.1% 91.0 ± 0.3% 99.7 ± 0.0% 97.6 ± 0.0% 99.2 ± 0.0% 96.8 ± 0.0%
0.50 88.4 ± 0.1% 78.7 ± 0.7% 96.7 ± 0.1% 92.0 ± 0.1% 99.8 ± 0.0% 97.8 ± 0.0% 99.2 ± 0.0% 96.9 ± 0.0%
Table 6: Table displaying values to form Figure 6. Accuracy after applying latent intervention for all tasks in test-ID and test-OOD for different levels of latent intervention of the color latent attribute for various numbers of colors. Altering the color hue randomly during training allows the model to gain up to 15% more OOD accuracy over the baseline.
Color Shape Size Material
Num. Colors Hidden Size Test-ID Test-OOD Test-ID Test-OOD Test-ID Test-OOD Test-ID Test-OOD
8 32 79.3 ± 0.1% 4.6 ± 0.6% 90.8 ± 0.8% 0.0 ± 0.0% 98.0 ± 0.0% 91.2 ± 1.0% 96.9 ± 0.0% 87.8 ± 0.9%
64 87.0 ± 1.6% 18.1 ± 1.9% 98.1 ± 0.4% 0.0 ± 0.0% 99.5 ± 0.0% 97.6 ± 0.0% 98.6 ± 0.0% 96.7 ± 0.1%
128 93.8 ± 0.1% 43.9 ± 0.7% 99.5 ± 0.0% 0.2 ± 0.0% 99.7 ± 0.0% 97.6 ± 0.0% 99.2 ± 0.0% 96.8 ± 0.0%
256 94.1 ± 0.1% 50.5 ± 1.1% 99.6 ± 0.0% 0.6 ± 0.1% 99.8 ± 0.0% 97.8 ± 0.0% 99.2 ± 0.0% 96.9 ± 0.0%
216 256 88.1 ± 0.1% 77.9 ± 0.4% 96.3 ± 0.0% 90.0 ± 0.4% 98.0 ± 0.0% 91.2 ± 1.0% 96.9 ± 0.0% 87.8 ± 0.9%
512 88.8 ± 0.0% 81.6 ± 0.2% 97.3 ± 0.0% 93.5 ± 0.1% 99.5 ± 0.0% 97.6 ± 0.0% 98.6 ± 0.0% 96.7 ± 0.1%
1024 89.1 ± 0.1% 82.2 ± 0.1% 97.6 ± 0.0% 92.4 ± 0.2% 99.7 ± 0.0% 97.6 ± 0.0% 99.2 ± 0.0% 96.8 ± 0.0%
Table 7: Table displaying values to form Figure 7. ID and OOD accuracy for models trained with different values for their hidden dimensions on a training set with 8 and 216 colors for all tasks.